OpenStudy (lordhelix8th):

Caley has a bag that contains 9 blue marbles and 13 yellow marbles. She selects a marble at random, and then, without replacing the first one, selects another marble at random. What is the probability that Caley selects a blue marble and then a yellow marble? Round your answer to the nearest percent. P(blue and yellow) ≈ %

OpenStudy (lordhelix8th):

@Compassionate @yuri_modoka

OpenStudy (lordhelix8th):

@HelpBlahBlahBlah

OpenStudy (lordhelix8th):

@MasterprojectHD @TheAsker2002 @toxicsugar22

OpenStudy (lordhelix8th):

@DarkMoonZ

OpenStudy (lordhelix8th):

@sleepyjess @sethRB26DETT

OpenStudy (lordhelix8th):

@amistre64

OpenStudy (lordhelix8th):

@Here_to_Help15

OpenStudy (lordhelix8th):

@Tootles143

OpenStudy (amistre64):

what the probability of picking a blue marble first?

OpenStudy (lordhelix8th):

@nincompoop

OpenStudy (lordhelix8th):

@JackofallTradez

OpenStudy (lordhelix8th):

it need to be shortend like 0.2

OpenStudy (amistre64):

The probabilty of the second pull is effected by the results of the first, so we have to work out the probability of dependant events. P(A|B) = P(AnB)/P(B) therefore, P(AnB) = P(B) P(A|B)

OpenStudy (amistre64):

its not that practical at this stage, but what it mean is: the probability of the first pull, times the adjustment for the second pull is what we are looking for

OpenStudy (amistre64):

P(blue first) = #blue, divided by # marbles P(yellow second) = #yellow remaining, divided by # of all marbles remaining.

OpenStudy (lordhelix8th):

so the answer is 25

OpenStudy (amistre64):

no how many blue marbles are there to start with? and how many marbles are there altogheter to start with?